20190515-Analyzer-Mtg

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> Analyzer_Meeting

previous meeting << >> following meeting

Logistic information

Meeting room at JLab: CC L210A (entrance is in the lobby across from the 2nd floor auditorium entrance)

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  419 045 306
 Bluejeans link: https://bluejeans.com/419045306

Agenda

Analyzer code updates

  1. Recent branch activity & pull requests
    • PK: two open pull requests that are being processed.
    • PK: there are a couple of issue that will probably abandon (#13, #17, #23, #32, )
    • PK: We need to setup the blinder and blinding string. The analyzer needs to understand if we are running longitudinal or transverse (#110)
    • PK: need BMod extractor (#108). RR: working on event level extractor
    • Issue97 : KP: we need to be able to have asymmetry and individual pre/post cuts. TY/CG: will take a look at it.
    • Issue77: PK: started looking at it and have some ideas how to proceed. Should have a draft in a week.
  2. regression analysis
    • RR: haplog 3763: we compared values and errors between LRB and postPan. The uncertainty is calculated differently between the two (taking errors before or after the regression). We decided to use the japan (before the regression) version — we also don’t use these errors in the correction (it’s just an indicator for the stability of the slopes).
    • RR: haplog 3768: compared with standard output fit and LRB. Consistency shows we are on track. Looking to figure out an appropriate number of events for a mini run.
  3. feedback
    • PK: will talk to Amali later today.
  4. Beam modulation
    • Updates to the bmwClient
      • PK: need to check the updated client (Ye created it). Will follow up with her.
    • Decoding and BMOD tree extraction
      • PK: for the decoding of the data, I’ll pick up the task. RR is working on the extractor class to separate the data.


Hardware/DAQ tasklist

  1. EPICS run-start and fast logger (and EPICS decoder to slow tree)
    • CG: waiting for run control DB implementation to collect all start of run scripts. BM: Scott Barcus agreed to implement the HRS start of run scripts.
  2. Timing: mostly understood, still small mysteries (LNE)
    • PK: will work on it after the injector tests next week. Should take about 1 afternoon. The CH and spectrometers will be started from the same signal in CH and the injector will have its own timing board.
  3. Synchronization-check system. w/help from Ed Jastrzembski
    • PK: will primary use the scalars. not sure how we can use the VQWKs directly. Cameron has setup a copy in the injector for the scalars and setup the decoding. We need to investigate further because we are not getting what we expect.
    • PK: for the digital sync check Ed has been working on it. CC: haven’t heard anything recently. Will check with him again.
  4. Clean up cabling, labels, fasteners - First need to examine all signals
    • CC: in progress replacing crate.
    • BM: we should check that there are screws on modules and make sure
  5. Clean up CRL -- eliminate so-called "dirty trick"
    • CC: that is Bob’s court. PK: we should look at the CH and injector CRLs once we finish up with the timing and beamMod client.
  6. Pedestal studies, aka cross-talk
    • KP: Caryn has been only working in the injector. She can see an issue in the laser hut (ppm level) but not much (ppb) upstairs. The DAQ is fine.
    • CG: who is taking care of the HRS crates and the CH. PK: we have a battery in the CH. TY: will take care of this. KP: it should have relevance how the detectors are connected.
  7. Green Monster rework: Adam.
    • AZ: the GM is going to be ready by the end of the day. RR: will work on it after that from Adam’s example.
  8. automatic production of run-dependent database
    • CG: still waiting on the computing center
  9. alarms like end-of-run voice
    • later
  10. e-logbook entries
    • later
  11. Document everything. Circuit diagrams, cable map, "how to".
    • CC: will make circuit diagrams.
  12. ADAQ and HDD system monitoring
    • CG: basic system working for hall A. Paul K and CG are on the emails. We should check if we need anything further.
    • PK: we should make sure we are using the parity disk properly. Start a clean up.
    • PK: working on tape volumes.
  13. Remote power supply cycling capability:
    • CC: asked Alex about this and he said there are some. Will investigate further.

ReSpin of APEX runs

runlist

  1. Aggregator Sakib:
    • SR: made aggregator plots for the 30uA runs and uploaded them to the haplog3729. Working on adding number of events to the aggregator. Run 1378 is odd. TY: For that run the RTP cell failed. KP: will investigate
  2. Pass 2: Paul
    • PK: we should have another pass to output the japan regression in the same location so that we can develop the aggregator collection tools (right now the regression output is coming from the postpan text files). KP: we should focus on a limited set.

Present

Wouter, Tao, Devi, Michael V., Paul K.

Adam, Bob, Cameron, Robert, Ciprian, Sakib

Excused